OpenStudy (anonymous):

Medals given Which of the following is in simplest form? twelve-twentieths nine thirty-seconds eleven sixty-sixths thirty-two fortieths

OpenStudy (anonymous):

\[ \frac{12}{20}~\frac{9}{32}~\frac{11}{66}~\frac{32}{40}\] Now can any of these be reduced?

OpenStudy (anonymous):